# Generative Ontology: When Structured Knowledge Learns to Create

### v1

@@ 9  Generalizability
9.1  Cross-Domain Application
Table 6 illustrates eight domains where Generative Ontology applies, each with its ontological components, generation outputs, and validation criteria.
@@ 9.1  Cross-Domain Application
The methodology follows a consistent six-step pattern regardless of domain: (1) domain analysis, identifying expert concepts, relationships, and validity criteria; (2) ontology encoding, mapping concepts to Pydantic models with validators; (3) generation signatures, defining typed input/output contracts with expert personas; (4) retrieval augmentation, indexing exemplars by ontology structure; (5) validation loop with structural, semantic, and adversarial checks plus iterative refinement; and (6) multi-agent specialization, assigning focused expertise and concerns per ontology subdomain.
9.2  When to Use Generative Ontology
The pattern is most valuable when four conditions hold: (1) the domain has structure: experts can articulate what makes outputs valid; (2) exemplars exist: a corpus of successful examples enables RAG and grounds generation; (3) validity is checkable: constraints can be encoded as validation functions; and (4) generation is valuable: the effort of building the ontology is justified by repeated generation needs. Game design exemplifies all four criteria: rich structure, thousands of documented games [1], clear validity criteria, and ongoing need for novel designs.
@@ 10.1  Key Insights
The Grammar Paradox.
Constraints enable creativity. A poet needs grammar not because grammar limits poetry but because grammar makes poetry possible. Our ontology schema, with its required fields, enum constraints, and validation rules, does not restrict the LLM’s creativity but provides the scaffolding within which meaningful creativity occurs.
Neither Alone Suffices.
Traditional ontology without generation is static, a map that cannot build cities. LLM generation without ontology is unconstrained, fluent outputs that collapse under scrutiny. The synthesis produces outputs that are both creative and valid, both novel and coherent.
The Whiteheadian Connection.
Generative Ontology offers a concrete instantiation of Whitehead’s metaphysics [16, 6]. The ontology provides eternal objects, abstract patterns existing as potentials. Generation produces actual occasions, concrete instantiations where form meets matter. Creativity is the movement from potentiality to actuality.
The Pattern Generalizes.
Although we developed Generative Ontology through games, the pattern applies wherever structured domains exist. Music, architecture, cooking, narrative, software design: each has its grammar, its ontology, its space of valid outputs.
10.2  Future Directions
This work opens several paths for exploration. Optimization: DSPy’s optimization capabilities [10] (BootstrapFewShot, MIPRO) can improve generation quality over time by learning from successful outputs; the ontology validation function we defined enables principled optimization via reward shaping. Interactive design: a human designer could collaborate with the system, providing high-level creative direction while the system handles structural completeness, with the ontology serving as a shared language between human and AI. Cross-domain transfer: ontologies developed for one domain may inform generation in related domains; game mechanisms and narrative structures share conceptual space, and the abstraction of Generative Ontology may enable cross-pollination. Ontology learning: we assumed the ontology is given, but AI systems could learn ontologies from corpora of successful examples; the induction of generative grammars from exemplars is a compelling research direction.

### v2

@@ 11  Generalizability
Table 10 illustrates eight domains where Generative Ontology applies. The pattern is most valuable when four conditions hold: (1) the domain has structure: experts can articulate what makes outputs valid; (2) exemplars exist: a corpus of successful examples enables RAG; (3) validity is checkable: constraints can be encoded as validation functions; and (4) generation is valuable: the effort of building the ontology is justified by repeated generation needs.
@@
12  Discussion
12.1  The Constraint Paradox in Data Form
Our ablation results provide empirical evidence for what we term the constraint paradox: adding constraints (C2, C3) eliminates structural errors without improving creative quality, yet the combination of constraints with architectural specialization (C4) produces the largest creative gains. Constraints alone suppress richness; constraints plus specialization enable it. This suggests a Constraint–Architecture Interaction Model: creative quality is a function of constraint expressiveness multiplied by architectural specialization, not simply additive.
12.2  The Gap as Feature
The one-point creative gap between generated and published designs (Study 2) is, paradoxically, encouraging. Published games represent decades of iterative playtesting, community feedback, and designer refinement, processes that no single generation pass can replicate. That the pipeline produces designs in the 7–8 range (“good, playable, interesting”) rather than the 8–9 range (“polished, elegant, replayable”) suggests the gap reflects not a fundamental limitation of the approach but the absence of iterative refinement. The dimensions where parity is already achieved (tension/drama and social interaction) may be those most amenable to specification-time design, while the gap dimensions (elegance and replayability) may require iterative playtesting to optimize.
12.3  The Whiteheadian Connection
Generative Ontology offers a concrete instantiation of Whitehead’s metaphysics [19, 6]. The ontology provides eternal objects, abstract patterns existing as potentials. Generation produces actual occasions, concrete instantiations where form meets matter. Creativity is the movement from potentiality to actuality. Our empirical results show this movement is not random: the constraints shape the space of possible actualizations, and architectural specialization navigates that space effectively.
12.4  Limitations
Several limitations qualify our findings. LLM-as-evaluator: While our ICC analysis demonstrates evaluator reliability, it does not establish validity: LLM scores may be consistently biased. Human calibration (comparing LLM scores to expert judgment) remains future work. No playtesting: Our evaluation assesses design specifications, not gameplay. Properties like elegance and replayability may manifest differently in play than in specification. Reference set: The 20 benchmark games, while spanning three complexity tiers, represent a convenience sample of well-known titles. A larger, more diverse reference set would strengthen generalizability. Single model: All generation and evaluation used Claude Sonnet 4. Cross-model evaluation would assess robustness. Temperature sensitivity: Evaluator ICC was measured at temperature 0; reliability at higher temperatures is untested.
@@ 13  Conclusion and Future Work
The Grammar Paradox. Constraints enable creativity. Our ablation study confirms this empirically: schema validation eliminates structural errors ($d=4.78$) while multi-agent specialization produces the largest creative gains (fun $d=1.12$, depth $d=1.59$).
Neither Alone Suffices. Traditional ontology without generation is static. LLM generation without ontology is unconstrained. The benchmark comparison shows that the synthesis produces designs that are structurally equivalent to published games, with a bounded creative gap that likely reflects the absence of iterative playtesting rather than fundamental architectural limitations.
Reliable Measurement Matters. Our test-retest analysis (ICC 0.578–0.989, median 0.873) demonstrates that LLM-based evaluation can be reliable when properly constrained. The cross-study validation differentiates “non-significant due to noise” from “non-significant due to genuine equivalence,” a methodological contribution applicable beyond game design.
The Pattern Generalizes. Although we developed Generative Ontology through games, the pattern applies wherever structured domains exist. Music, architecture, cooking, narrative, software design: each has its grammar, its ontology, its space of valid outputs.
Future directions include human calibration (validating LLM scores against expert judgment), iterative design loops (multiple generation–evaluation–refinement cycles to close the creative gap), cross-model evaluation (assessing robustness across LLMs), and DSPy optimization (using the ontology validation function as a reward signal for automatic prompt tuning).
@@
Appendix A Theme Prompts
Table 11 lists the ten standardized theme prompts used in the ablation study. Each prompt was run under all four experimental conditions with three replications, yielding 120 total designs.
Table 11: Standardized theme prompts for the ablation study.
# Constraints Theme
1 2–4p, competitive, medium, 45–60min Bioluminescent fungi competing in a cave ecosystem
2 2–4p, competitive, med-heavy, 60–90min Victorian-era detective solving mysteries through deduction
3 3–5p, competitive, med-heavy, 60–90min Space traders negotiating routes between planets
4 2–4p, competitive, light, 30–45min Medieval knights jousting in a tournament
5 2–4p, cooperative, medium, 45–60min Deep sea divers recovering treasure from a sunken ship
6 2–4p, competitive, medium, 45–60min Rival chefs competing in a cooking competition
7 2–5p, competitive, med-heavy, 90–120min Ancient civilizations building wonders of the world
8 2–4p, cooperative, medium, 30–45min Escape room where the walls are closing in
9 2–4p, competitive, medium, 45–60min Fairy tale creatures vying for control of an enchanted forest
10 2–4p, cooperative, heavy, 90–120min Time travelers preventing paradoxes across historical eras
Appendix B Reference Games
Table 12 lists the 20 published board games used as the reference set in Study 2. Games were selected to span three complexity tiers (light, medium, heavy) and represent widely recognized titles with established design reputations. Each game was reverse-engineered into a complete GameGrammar ontology (25 required fields plus 3 balance parameters) and evaluated through the identical Design Coach pipeline used for generated designs.
Table 12: Reference games for the benchmark study (Study 2).
# Game Designer(s) Tier
1 Azul Michael Kiesling Light
2 Codenames Vlaada Chvátil Light
3 Kingdomino Bruno Cathala Light
4 Love Letter Seiji Kanai Light
5 Splendor Marc André Light
6 7 Wonders Antoine Bauza Medium
7 Catan Klaus Teuber Medium
8 Concordia Mac Gerdts Medium
9 Dune: Imperium Paul Dennen Medium
10 Everdell James A. Wilson Medium
11 Root Cole Wehrle Medium
12 Spirit Island R. Eric Reuss Medium
13 Terraforming Mars Jacob Fryxelius Medium
14 Ticket to Ride Alan R. Moon Medium
15 Wingspan Elizabeth Hargrave Medium
16 Brass: Birmingham Gavan Brown & Matt Tolman Heavy
17 Gaia Project Jens Drögemüller & Helge Ostertag Heavy
18 Gloomhaven Isaac Childres Heavy
19 Through the Ages Vlácil Chvátil Heavy
20 Twilight Imperium 4E Dane Beltrami & Corey Konieczka Heavy